OpenStudy (quickstudent):

Is this correct? question: As the marketing designer for Burger King, you are given a 4 in by 6 in logo and have been told to create an advertisement with sides six times larger. What is the area of your advertisement (in feet)? [12 inches = 1 ft] my answer: 4/6 * 6/6 = 24/36 The dimentions of the larger ad is 24 in by 36 in. 24 * 36 = 864 864/12 = 72 The area of the ad is 72 square feet.

OpenStudy (quickstudent):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

I agree to the point where you said that the new dimensions are 24 in by 36 inches.

OpenStudy (joannablackwelder):

From there, I would convert these lengths into feet before finding the area.

OpenStudy (joannablackwelder):

We can convert once we find the area, but it is a little tricky.

OpenStudy (joannablackwelder):

So, how many feet is 24 inches?

OpenStudy (quickstudent):

Oh, ok, I get it, here's my correction: 4/6 * 6/6 = 24/36 The dimentions of the larger ad is 24 in by 36 in. 24/12 = 2 36/12 = 3 2 * 3 = 6 The area of the ad is 6 square feet.